OpenStudy (anonymous):

USE THE SUBSTITUTION METHOD TO SOLVE 3M+N=20 M-8N=15

OpenStudy (anonymous):

to use substitution, we need to solve for one of the variables. Im choosing to use the second equation and solve for M because it looks easy, so we obtain: \[M-8N = 15 \Rightarrow M = 8N+15\] Now Im going to substitute that value of M into the first equation; everywhere there is an M in the first equation I will replace it with 8N+15: \[3M+N+20 \Rightarrow 3(8N+15)+N=20 \Rightarrow 24N+45+N=20\] \[25N=-25 \Rightarrow N=-1\] now that I have N = -1, i can put that back into the first equation I had: \[M=8N+15 \Rightarrow M=8(-1)+15 \Rightarrow M = -8+15 \Rightarrow M = 7\]